OpenStudy (anonymous):

Find the value of term a^14 in the sequence. 6,5,4,3,2

OpenStudy (anonymous):

a^14? Or do you mean a sub 14...the 14th term?

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

The sequence is: a + (n-1)d where a = first term, n = nth term, d = difference. So in 6,5,4,... the difference is 5-6 = -1 and the first term, a , is 6. so the sequence is 6 + (n-1)(-1) which is 6 - n + 1 = 7-n So the sequence is represented by 7-n. So if you want the 14th term, substitute 14 for n, you get 7-14 = -7. The 14th term is -7.

OpenStudy (anonymous):

I understand alittle bit

OpenStudy (anonymous):

In an arithmetic sequence, you know how to find the common difference, d, take the second term - first term, or take third term - second term...and in this case you get -1. You know the first term is 6. So just substitute into the formula a+(n-1)d

OpenStudy (anonymous):

oh okay im taking math on the computer it's harder without a teacher

OpenStudy (anonymous):

Understood.